Two planes depart from the same place at 2pm. One plane flies south at a speed of 376km/hr, and the other flies west at a speed of 648km/hr. How far apart are the airplanes at 5:30PM?
Answer by mananth(16949)

Two planes depart from the same place at 2pm. One plane flies south at a speed of 376km/hr, and the other flies west at a speed of 648km/hr. How far apart are the airplanes at 5:30PM?
distance traveled by first plane to south in 3.5 hours= 376*3.5=1316
distance traveled by first plane to west in 3.5 hours= 648*3.5=2268
they are traveling at right angles
so apply pythagoras theorem
for finding the hypotenuse = distance
(1316)^2 +(2268)^2 = 1731856 + 5225796 = (6875680)^2 = 2622.15